# On the cross-correlation properties of large-size families of Costas arrays

### v1

@@ V Conclusion and future work
In this paper, we investigated several large-size families of Costas arrays and extended arrays, and gave certain bound on the maximal cross-correlation of these families. More precisely, we gave an upper bound on the maximal cross-correlation of the family $\mathcal{W}^{el}_{p}$ for arbitrary $u$ and $v$. Additionally, we provided an upper bound for the power permutation family $\mathcal{P}_{p}$ in the case $u=0,v\neq 0$, utilizing a result on trinomials over finite fields [18]. Furthermore, through combinatorial arguments, we obtained the first nontrivial upper bound on the maximal cross-correlation of the family $\mathcal{PW}_{p}$ for arbitrary $u$ and $v=0$, thereby answering an open problem posed by Ardalani [10, p. 117]. A possible future research direction can be devoted to tighten the upper bounds on the existing families of Costas arrays, and present new large-size families.

### v2

@@ V Conclusion and future work
In this paper, we investigated several large-size families of Costas arrays and extended arrays, and gave certain bounds on the maximal cross-correlation of these families. More precisely, we gave an upper bound on the maximal cross-correlation of the family $\mathcal{W}^{el}_{p}$ for arbitrary $u$ and $v$. Additionally, we provided an upper bound for the power permutation family $\mathcal{P}_{p}$ in the case $u=0,v\neq 0$, utilizing a result on trinomials over finite fields [KO2017trinomials]. Furthermore, through combinatorial arguments, we obtained the first nontrivial upper bound on the maximal cross-correlation of the family $\mathcal{PW}_{p}$ for arbitrary $u$ and $v=0$, thereby answering an open problem posed by Ardalani [Ardalani2023Contributions, p. 117]. A possible future research direction can be devoted to tighten the upper bounds on the existing families of Costas arrays, and present new large-size families.